Hi,
I have been working this problem and I can not get the proper answer. Please help!
Whitney invested $38,000, part at 2% and part at 12%. If the total interest at the end of the year is $2,760, how much did she invest at 2%?
Thank you
Sherri

Answer by stanbon(48516)

Let amount invested at 2% be "x"
Then amount invested at 12% is "38000-x"
Interest at 2% is 0.02x
Interest at 12% is 0.12(38000-x)= 4560-o.12x
EQUATION:
Int. at 2% + Int. at 12% = $2760
0.02x + 4560-0.12x = 2760
2x + 456000- 12x = 276000
-10x = -180000
x = $18,000 (amount invested at 2%)
38000-x = $20000 (amount invested at 12%)
Cheers,
Stan H.